OpenStudy (anonymous):

Find the solution of this system of equations algebraically. (A) y = -x (B) y = x3 + 3x2 + 2x (0, 0) (0, -1) and (0, -2) (0, 0), (-1, 1) and (-2, 2) (-1, 1) and (-3, 3) It has no solution.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

If we apply the elimination method, we can write: \[\Large \begin{gathered} - x = y = {x^3} + 3{x^2} + 2x \hfill \\ {x^3} + 3{x^2} + 3x = 0 \hfill \\ x\left( {{x^2} + 3x + 3} \right) = 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

It also has complex roots

OpenStudy (michele_laino):

yes! Nevertheless we have to understand if there are other solutions or there are not

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so what is the right option?

OpenStudy (anonymous):

ugh the one with no solution?

OpenStudy (michele_laino):

we have only one real solution

OpenStudy (michele_laino):

namely x=0, y=0

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks!

OpenStudy (michele_laino):

thanks! :)